OpenStudy (anonymous):

Resolve into partial fractions 1/[ (x^2 +1 )^2 (x-3) ]

OpenStudy (jiteshmeghwal9):

http://answers.yahoo.com/question/index?qid=20100811085245AAwrqXD look @ this @Yahoo! :)

OpenStudy (anonymous):

Is it : \[1 = \frac{Ax + D}{(x^2 + 1)} + \frac{Bx + E}{(x^2 + 1)^2} + \frac{C}{(x-3)}\] I am not sure..

OpenStudy (anonymous):

yes it is !!!

OpenStudy (anonymous):

@lgbasallote help here..

OpenStudy (jiteshmeghwal9):

do this works @Yahoo! ?????

OpenStudy (anonymous):

If it right then half of our work is finished..

OpenStudy (lgbasallote):

No.

OpenStudy (anonymous):

you can check it here your decomposition is correct ! http://www.wolframalpha.com/input/?i=Partial+fractions1%2F%5B+%28x%5E2+%2B1+%29%5E2+%28x-3%29+%5D

OpenStudy (anonymous):

Sorry there will come denominator also..

OpenStudy (anonymous):

No matters.. They will cancel out after taking LCM..

OpenStudy (anonymous):

Now let us find A, B , C D and E..

OpenStudy (anonymous):

@waterineyes i have the answer key

OpenStudy (anonymous):

Keep it with yourself..

OpenStudy (anonymous):

i think u r correct plzz procede

OpenStudy (anonymous):

After taking LCM we get: \[1 = (Ax + D)(x^2 + 1)(x-3) + (Bx +E)(x^2 + 1)(x-3) + C(x^2 + 1)^2\]

OpenStudy (anonymous):

now compare coefficients..: firstly for \(x^4\).. A + B + C = 0 Put x = 3: C(10) = 1 C = 1/10

OpenStudy (anonymous):

*100.

OpenStudy (anonymous):

\[C = \frac{1}{100}\]

OpenStudy (anonymous):

Sorry I have done wrong..

OpenStudy (anonymous):

\[1 = (Ax + D)(x^2 + 1)(x-3) + (Bx +E)(x-3) + C(x^2 + 1)^2\]

OpenStudy (anonymous):

From here when putting x = 3 we get: \[C =\frac{1}{100}\]

OpenStudy (anonymous):

Compare the coefficients of \(x^4 \) here: A + C = 0 From here: \[A = \frac{-1}{100}\]

OpenStudy (anonymous):

Sorry: \[D = \frac{-3}{10}\]

OpenStudy (anonymous):

Now you have to find B..

OpenStudy (anonymous):

Yes I am..

OpenStudy (anonymous):

Have to do once again.. A C and D is right..

OpenStudy (anonymous):

B and E must be checked once again..

OpenStudy (anonymous):

No D is not right..

OpenStudy (anonymous):

Yeah I got for D.. Check the coefficient of \(x^3\) there : \[-3A + D = 0\] \[D = 3A \implies D = \frac{-3}{100}\]

OpenStudy (anonymous):

Now we got A, C and D..

OpenStudy (anonymous):

Compare the constant terms here: \[-3D -3E + C = 1 \implies -3E = 1- \frac{1}{100} - \frac{9}{100} \implies \frac{90}{100} \implies E = \frac{-3}{10}\]

OpenStudy (anonymous):

Compare the coefficients of \(x\) here: \[-3A + D - 3B + E = 0\] \[3B = \frac{3}{100} - \frac{3}{100} - \frac{3}{10} \implies 3B = \frac{-3}{10} \implies B = \frac{-1}{10}\]

OpenStudy (anonymous):

So you have all the values now: \[A = -\frac{1}{100}, D = -\frac{3}{100}, B = - \frac{1}{10}, E = -\frac{3}{10} \; and \; C = \frac{1}{100}\]

OpenStudy (anonymous):

So now you can rearrange your expression to: \[\implies \color{blue}{\frac{-x -3}{100(x^2 + 1)} + \frac{-x -3}{10(x^2 + 1)^2} + \frac{1}{100(x-3)}}\] Getting @Yahoo!

OpenStudy (anonymous):

Yup .... thxx

OpenStudy (anonymous):

Welcome dear..